OpenStudy (anonymous):

Differentiate the functions. Assume that C is constant. g(x)=7^x

OpenStudy (anonymous):

Derivative of a^x (where a is a constant) is a^x (ln a) So the derivative of 7^x is 7^x (ln 7)

OpenStudy (anonymous):

@Easyaspi314 So is this is the same as (ln7)7^x right? I don't understand why we have to take ln7?

OpenStudy (anonymous):

Yes it is the same.

OpenStudy (anonymous):

Keep in mind, that the derivative of e^x would follow the same rule, it will be e^x (ln e)..but ln e = 1 by definition, so the derivative of e^x is just e^x. It can be shown that the derivative of a^x is a^x (ln a).

OpenStudy (anonymous):

If you want a proof, you can check your text or I can help you with it/

OpenStudy (anonymous):

@EasyasPi314 My text does not do much of an explanation actually.

OpenStudy (anonymous):

Did I answer your question, or do you need more help?

OpenStudy (anonymous):

It would be helpful if dont mind explaining.

OpenStudy (anonymous):

Consider y= a^x. Take the log of both sides, you get ln y = x ln a Now, take the derivative of both sides, you get 1/y (y ') = ln a Now multiply both sides by y, you get y ' = y(ln a) Therefore, y', the derivative of y, is y(ln a) OR (a^x)(lna) So you clearly see the poroof that if y = a^x Then: y ' = a^x (ln a)

OpenStudy (anonymous):

Makes sense?????

OpenStudy (anonymous):

yes this does makes sense, thanks @ Easaspi314

OpenStudy (anonymous):

welcome.